# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v2

@@ 4 Defining Reward Hacking
Despite the difficulties in formalizing reward hacking, we argue that both of our desiderata for a definition can be met by defining reward hacking with respect to a base policy. We show how this allows for both a precise definition of proxy rewards and a natural threshold for when reward hacking is occurring.
Characterizing proxy rewards.  To find a realistic definition of a good proxy reward, consider the process by which system designers create proxies. If they are hand-specified, we argue that designers usually imagine a distribution of behavior and design a reward that captures the objective under that distribution. For example, we study a traffic control simulator (Lopez et al., 2018; Vinitsky et al., 2018; Wu et al., 2022; Pan et al., 2022) where a small number of autonomous cars help to regulate traffic among a larger population of human drivers. In this case, a designer might choose the average speed of all vehicles as a proxy for improving traffic flow; this is a reasonable proxy because under the distribution of typical human driving behavior, higher speeds are associated with better traffic.
@@
$\qquad\qquad\quad\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\left(\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}\right)\left(\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\right)\right]=r,$
@@
$\textstyle\text{where}\quad\sigma_{\tilde{R}}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})\big)^{2}\right]\quad\text{and}\quad\sigma_{R}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(R(s,a)-J({\pi_{\text{base}}},R)\big)^{2}\right]$
are the standard deviations of proxy and true rewards under the base policy.
@@
This definition intuitively captures cases like the traffic environment, where we define the true reward function as the negative total commute time for all vehicles. Letting $\pi_{\text{base}}$} be an autonomous vehicle policy based on a common model of human driving behavior, we plot the distribution of true and proxy rewards in the top-left corner of Figure 2 and find that they are correlated: higher average speed tends to occur with lower commute times, and vice versa. This validates that average speed is a correlated proxy reward according to Definition 4.1.
Definition 4.1 also avoids too strongly constraining proxy rewards, unlike past formalisms. The proxy and true reward functions can diverge arbitrarily at some state-action pairs, as long as those state-action pairs have low or zero occupancy measure under the base policy.
@@
Finally, Definition 4.1 captures learned proxy rewards in addition to hand-specified ones. Reward learning begins with collecting annotations—for example, preference comparisons (as in RLHF) or scalar feedback—for states and actions collected from a rollout policy—for example, the SFT model in RLHF. Then, a reward function is estimated via supervised learning over this dataset with an appropriate loss function. If the estimated reward function generalizes well in-distribution, then we should expect the true and learned rewards to be well correlated under states and actions sampled from the rollout policy, satisfying Definition 4.1.
@@
Choosing a threshold for reward hacking.  As discussed in Section 3, the other difficulty in defining reward hacking is specifying a threshold of true reward below which performance is poor enough to be considered hacking. If we are already characterizing proxy rewards with respect to a base policy, then it makes sense to also use this base policy as a baseline for evaluating a policy that optimizes the proxy. If optimizing the proxy leads to worse true reward than the base policy achieves, then the system designer may as well simply use the base policy. On the other hand, if optimizing the proxy reward improves true reward over the base policy, then we can argue that optimizing the proxy succeeded.
Definition 4.2 (Reward hacking).
@@
Suppose a $\tilde{R}$ is an $r$-correlated proxy with respect to $\pi_{\text{base}}$} (Definition 4.1). Then we say reward hacking occurs when an optimal policy for $\tilde{R}$ has lower true reward than the base policy $\pi_{\text{base}}$}:
@@
For example, in the traffic control environment, optimizing for the proxy of average speed leads to the autonomous vehicles blocking a highway on-ramp. This decreases the speed of the cars on the ramp to zero, but allows cars on the highway to move at high speeds, overall increasing the proxy reward. However, the true reward becomes extremely low, since commute times for the cars on the on-ramp are aribitrarily long. This constitutes reward hacking under Definition 4.2 since the true reward of the proxy-optimized policy is lower than that achieved by typical human driving.
Verifying our definition experimentally.  To test whether Definition 4.2 accurately captures intuitive cases of reward hacking, we consider four realistic enviroments and an illustrative gridworld. We consider the aforementioned traffic simulator and two other environments originally studied by Pan et al. (2022) as examples of reward hacking. SimGlucose (Man et al., 2014) is based on an FDA-approved simulator of Type 1 Diabetes patients in which a policy monitors glucose levels and administers insulin; the true reward captures patient health while the proxy reward prioritizes reducing the monetary cost of insulin. PandemicSimulator (Kompella et al., 2020) uses the SEIR model (Mwalili et al., 2020) to simulate the COVID-19 pandemic among a population; the policy controls the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward omits the political cost of certain decisions.
We also study RLHF, in which LLMs are optimized based on a reward function learned from human preferences. Following Gao et al. (2022) and Coste et al. (2024), we use a large (more robust) reward model as the true reward function and a smaller (less robust) one as the proxy. Finally, as an interpretable and illustrative example, we include the tomato-watering gridworld from Leike et al. (2017). In this environment, a robot that waters tomatoes receives true reward depending on how many plants are watered; however, the proxy additionally rewards the robot for standing in a sprinkler where it appears all the tomatoes are watered. Besides the gridworld, all of our environments reflect complex, realistic tasks with large or infinite state state spaces. To test our definition of reward hacking, for each of the five environments we construct a natural base policy. See Appendix C for more details about the environments and base policies.
@@
Figure 3: Unlike RLHF, which regularizes action distribution (AD) divergence to prevent reward hacking, our results suggest regularizing using occupancy measure (OM) divergence. These plots of the glucose monitoring environment show the typical ADs and OMs of two policies. $\pi$ is close to $\pi_{\text{base}}$} in AD; it is more likely to give slightly less insulin because the proxy reward penalizes monetary cost. However, this leads to a vastly different OM, with typical glucose levels for $\pi$ far outside the healthy range (dotted lines). Thus, regularizing ADs to be close to $\pi_{\text{base}}$} is not enough to prevent reward hacking; instead, divergence between the OMs better captures the reward hacking behavior.
@@ 5 Mitigating Reward Hacking with Occupancy Measure Regularization
We now discuss how our new definition of reward hacking can lead to better methods for preventing it. Ideally, we would like to be able to optimize a proxy reward function and have it translate into an improvement in true reward over the base policy. To understand how this might be possible, consider again the traffic environment. The reward hacking policy exhibits behavior which is very unlikely under the base policy: human drivers hardly ever stop on on-ramps and refuse to move. That is, optimizing the proxy reward leads to out-of-distribution states and actions where the correlation that made the proxy good in the first place breaks down. Visually, this can be seen in the second row of Figure 2: the reward hacking policy usually finds state-action pairs with high proxy reward and low true reward that were very unlikely to be reached by the base policy.
Thus, to prevent reward hacking, one solution could be to optimize the proxy reward while avoiding states that are unlikely under the base policy. The following theorem formalizes this idea.
@@
See Appendix A for the proof. Equation (4) gives a lower bound on how much the policy $\pi$ improves over the base policy $\pi_{\text{base}}$} on the true reward, normalized by the standard deviation of the true reward under the base policy. This is exactly we would like to maximize, but we can’t optimize the unknown true reward directly, so it makes sense to instead optimize the right hand side of (4). The RHS consists of two terms. The first is the improvement of the policy $\pi$ over the base policy under the proxy reward, again normalized; optimizing this term alone often leads to reward hacking. However, the second term penalizes the divergence of $\pi$’s occupancy measure from that of $\pi_{\text{base}}$}. By incentivizing $\pi$ to achieve high proxy reward but also stay close to the base policy, $\pi$ can achieve high true reward.
Scaling the RHS of (4) and removing terms that are constant in $\pi$ suggests using the following regularized policy optimization objective to avoid reward hacking:
@@
Comparison to KL regularization in RLHF.  Regularization to a base policy is already widely used to prevent reward hacking in RLHF (Stiennon et al., 2020; Bai et al., 2022). Specifically, a KL penalty is applied between the distributions of responses generated by the optimized policy and the SFT policy. In our setting, we can write this as
@@
Our regularized objective in (5) differs in two main ways from the KL regularization used in RLHF. First, our results suggest optimizing the occupancy measure (OM) divergence between policies, whereas RLHF uses the action distribution (AD) divergence. Second, Theorem 5.1 applies to $\chi^{2}$ divergence, while RLHF uses KL divergence. In the remainder of this section, we explore intuitively why OM divergence is preferable to AD divergence, and why $\chi^{2}$ divergence is superior to KL divergence. Then, in Section 6, we empirically explore applying different types of regularization to prevent reward hacking in five environments.
@@
Figure 4: Our theory also suggests preventing reward hacking by regularizing with $\chi^{2}$ divergence instead of KL divergence like prior work. This plot illustrates how $\chi^{2}$ regularization is more effective at preventing reward hacking in RLHF. Both divergences can be written as the expectation of a function $g\big(\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big)\big)$ which penalizes state-action pairs more the further the log-ratio is from zero. The $g(\cdot)$ associated with KL divergence only increases slowly for large log-ratios, so policies trained with KL divergence may produce nonsensical text. In contrast, the $g(\cdot)$ for $\chi^{2}$ divergence increases exponentially, better constraining the LLM to produce text similar to the SFT policy.
Occupancy measure vs. action distribution regularization.  While AD regularization works well for RLHF, this may be because RLHF is essentially a contextual bandit problem, meaning that $\gamma=0$; in this case, OM and AD divergence are equivalent (see Appendix A.3). However, in other cases, AD divergence may not suffice to prevent reward hacking behavior. This is because, in longer-horizon environments, a small change in action distribution at a single state can lead to a much higher probability of reaching undesirable states. Figure 3 shows an example of this in the glucose monitoring environment: policies that are close in action distribution regularization produce vastly different patient glucose levels. Occupancy measure regularization avoids this issue by directly preventing the distribution of glucose levels from differing too much from the base policy.
$\chi^{2}$ vs. KL divergence.  Compared to KL divergence, $\chi^{2}$ divergence may be superior for preventing reward hacking because it more strongly penalizes out-of-distribution state-action pairs. To illustrate this, we can write both divergences as expectations over functions of the log-ratio of the occupancy measures, which we denote $d(s,a)$:
@@
Table 1: We compare using various types of regularization to prevent reward hacking in the five environments form Figure 2. The median true reward and standard deviation across 5 random seeds is shown for the best regularization coefficient for each type of regularization. The bottom rows show results for the baselines: the base policy $\pi_{\text{base}}$}, a policy trained on the proxy reward without regularization (exhibiting reward hacking), and a policy trained on the true reward function (impossible in practice, but included as an upper bound on performance). We find that occupancy measure regularization consistently improves on action distribution regularization, and that $\chi^{2}$ divergence is often superior to KL divergence.
@@ 6 Experiments
We now show that our theoretical results—which suggest $\chi^{2}$ occupancy measure regularization can prevent reward hacking—lead to empirical success in realistic environments.
Practical occupancy measure regularization.  Occupancy measure regularization is more difficult to implement in practice compared to action distribution regularization. While AD regularization can be added as a loss term to deep RL algorithms like proximal policy optimization (PPO), OM divergences cannot be calculated in closed form. Instead, we follow several previous works (e.g., Ho & Ermon 2016; Kang et al. 2018) and use a discriminator network to approximate OM divergences. Specificaly, in Appendix B, we show the objective in (5) can be optimized via policy gradient with an adjusted reward function that depends on a discriminator $\hat{d}_{\phi}$:
@@
$\text{and}\qquad\phi$ $=\arg\min_{\phi}\;\mathbb{E}_{\mu_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi}(s,a)})\,\Big]+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi}(s,a)})\,\Big].$ (8)
That is, optimizing a discriminator network $\hat{d}_{\phi}$ to minimize the given loss can be used to estimate and optimize $\chi^{2}$ OM divergence. We alternately train $\hat{d}_{\phi}$ via gradient descent on (8) and the policy $\pi$ via PPO based on the adjusted reward. We call this algorithm Occupancy-Regularized Policy Optimization (ORPO). See Appendix B for a full derivation of the approximations used in ORPO and Algorithm 1 for a formal description. We use a similar strategy for regularizing based on OM KL divergence; see the appendix for details.
@@
However, comparing the types of regularization based only on the best coefficient is unrealistic, since in practice designers must choose a coefficient $\lambda$ without access to the true reward. In Figure 5, we show the true reward achieved when training with each type of regularization across a range of $\lambda$ values. In addition to performing best with an optimal coefficient, $\chi^{2}$ and OM regularization seem to also perform well over a larger range of coefficients compared to AD regularization.
In Appendix E, we present the full results of our experiments as well as ablations of ORPO.

### v3

@@ 4 Defining Reward Hacking
Despite the difficulties in formalizing reward hacking, we argue that both of our desiderata for a definition can be met by defining reward hacking with respect to a base policy. This approach enables both a precise definition of proxy rewards and a natural true reward threshold for determining when reward hacking is occurring.
Characterizing proxy rewards.  To find a realistic definition of a good proxy reward, consider the process by which system designers create proxies. If they are hand-specified, we argue that designers usually imagine a distribution of behavior and design a reward that captures the objective under that distribution. For example, we study a traffic control simulator (Lopez et al., 2018; Vinitsky et al., 2018; Wu et al., 2022; Pan et al., 2022) where a small number of autonomous cars help to regulate traffic among a larger population of human drivers. In this case, a designer might choose average vehicle speed as a proxy for traffic flow; this is a reasonable proxy because under the distribution of typical human driving behavior, higher speeds are associated with better traffic conditions.
@@
$\textstyle\qquad\qquad\quad\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\left(\frac{\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}\right)\left(\frac{R(s,a)-J({\pi_{\text{base}}},R)}{\sigma_{R}}\right)\right]=r,$
@@
$\textstyle\text{where}\quad\sigma_{\tilde{R}}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(\tilde{R}(s,a)-J({\pi_{\text{base}}},\tilde{R})\big)^{2}\right]\quad\text{and}\quad\sigma_{R}^{2}=\mathbb{E}_{\mu_{\pi_{\text{base}}}}\left[\big(R(s,a)-J({\pi_{\text{base}}},R)\big)^{2}\right]$
are the variances of proxy and true rewards under the base policy.
@@
To consider an application of our definition, let’s return to the example of our traffic environment, where we define the true reward function as the negative total commute time for all vehicles. Using an autonomous vehicle policy based on a common model of human driving behavior as $\pi_{\text{base}}$}, we plot the distribution of true and proxy rewards in the top-left corner of Figure 2. The clear correlation that we observe between the two reward functions—higher average speed corresponding to lower commute times and vice versa—confirms that average speed is a correlated proxy reward according to Definition 4.1. This demonstrates that our formal definition successfully captures intuitive real-world proxy rewards.
@@
Definition 4.1 offers key advantages over past formalisms. First, it avoids too strongly constraining proxy rewards by allowing the proxy and true reward functions to diverge arbitrarily at some state-action pairs when those pairs have low or zero occupancy measure under the base policy. Second, it encompasses both learned proxy rewards and hand-specified ones. In reward learning, we collect annotations (such as preference comparisons or scalar feedback) for states and actions collected from a rollout policy (such as the SFT model in RLHF). Then, a reward function is estimated via supervised learning over this dataset with an appropriate loss function. If the estimated reward function generalizes well in-distribution, then we should expect the true and learned rewards to be well-correlated under states and actions sampled from the rollout policy, satisfying Definition 4.1 (see Lemma A.8).
@@
Choosing a threshold for reward hacking.  As discussed in Section 3, defining reward hacking requires specifying a threshold of true reward below which policy behavior is considered poor enough to be classified as reward hacking. Given that we already characterize proxy rewards with respect to a base policy, this same base policy provides a natural baseline for evaluating a policy that optimizes the proxy: if optimizing the proxy leads to worse true reward than the base policy achieves, then the system designer may be better off simply using the base policy.
Definition 4.2 (Hackable proxy reward).
@@
Suppose $\tilde{R}$ is an $r$-correlated proxy with respect to $\pi_{\text{base}}$} (Definition 4.1). Then we say reward hacking occurs when a policy $\pi$ optimized for $\tilde{R}$ has lower true reward than the base policy $\pi_{\text{base}}$}, i.e., when $J(\pi,R)<J({\pi_{\text{base}}},R)$. If an optimal policy for $\tilde{R}$ exhibits reward hacking, then we say that the proxy reward $\tilde{R}$ is hackable:
@@
For example, in the traffic control environment, optimizing average speed as a proxy reward leads to autonomous vehicles blocking highway on-ramps, a strategy that increases overall average speed by allowing highway traffic to move freely while forcing on-ramp cars to stop completely. However, the true reward becomes extremely low because commute times for the cars on the on-ramp are now arbitrarily long. This policy is thus reward hacking under Definition 4.2 since the true reward that it accumulates is lower than that achieved by typical human driving.
Verifying our definition experimentally.  To test whether Definition 4.2 accurately captures intuitive cases of reward hacking in practice, we consider four realistic environments and an illustrative gridworld. In addition to the aforementioned traffic simulator, we consider a pandemic mitigation and glucose monitoring environments, all of which were originally studied by Pan et al. (2022) as examples of reward hacking. PandemicSimulator (Kompella et al., 2020) uses a specialized SEIR (susceptible, exposed, infected, recovered) infection model to simulate the COVID-19 pandemic among a population; the policy controls the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward omits the political cost of certain decisions. SimGlucose (Man et al., 2014) is based on an FDA-approved simulator of Type 1 Diabetes patients in which a policy monitors glucose levels and administers insulin. The true reward captures patient health while the proxy reward prioritizes reducing the monetary cost of insulin.
We also study RLHF, in which LLMs are optimized based on a reward function learned from human preferences. Following Gao et al. (2022) and Coste et al. (2024), we use a large (more robust) reward model as the true reward function and a smaller (less robust) one as the proxy. Finally, as an interpretable and illustrative example, we include the tomato-watering gridworld from Leike et al. (2017). In this environment, a robot that waters tomatoes receives true reward depending on how many plants are watered; however, the proxy additionally rewards the robot for standing in a sprinkler where it appears all the tomatoes are watered. Besides the gridworld, all of our environments reflect complex, realistic tasks with large or infinite state state spaces. To test our definition of reward hacking, we construct a natural base policy for each of the five environments. See Appendix C for more details about the environments and base policies.
@@
Figure 3: Unlike RLHF, which attempts to prevent reward hacking by regularizing action distribution (AD) divergence from the base policy, our results suggest regularizing using occupancy measure (OM) divergence is more effective. These plots of the glucose monitoring environment show the typical ADs and OMs of two policies. $\pi$ is close to $\pi_{\text{base}}$} in AD; it is more likely to give slightly less insulin due to the proxy reward’s monetary cost penalty. However, $\pi$’s optimization of the proxy also leads to a vastly different OM with typical glucose levels far outside the healthy range (dotted lines). Thus, regularizing ADs to be close to $\pi_{\text{base}}$} is not enough to prevent reward hacking. Instead, divergence between the OMs better captures the reward hacking behavior.
@@ 5 Mitigating Reward Hacking with Occupancy Measure Regularization
We now discuss how our new definition of reward hacking motivates better methods for preventing it. Ideally, we would like to be able to optimize a proxy reward function and have it translate into an improvement in true reward over the base policy. To understand how this might be possible, consider again the traffic environment. The reward hacking policy exhibits behavior which is very unlikely under the base policy: human drivers hardly ever stop indefinitely on on-ramps. That is, optimizing the proxy reward leads to out-of-distribution states and actions where the correlation that made the proxy good in the first place breaks down. Visually, this can be seen in the second row of Figure 2: the reward hacking policy usually finds state-action pairs with high proxy reward and low true reward that were very unlikely to be reached by the base policy.
Thus, one solution to prevent reward hacking could be to optimize the proxy reward while avoiding states that are unlikely under the base policy. The following theorem formalizes this idea.
@@
See Appendix A for the proof. Equation (10) gives a lower bound on how much the policy $\pi$ improves over the base policy $\pi_{\text{base}}$} in terms of the true reward, normalized by the standard deviation of the true reward under the base policy. Although this is exactly what we would like to maximize, we must instead optimize the right-hand side (RHS) of (10) since we cannot access the true reward directly. The RHS consists of two terms: one measuring the normalized improvement of the policy $\pi$’s proxy reward over the base policy and the other penalizing the divergence of $\pi$’s occupancy measure from that of $\pi_{\text{base}}$}. Optimizing the first term alone often leads to reward hacking, but by both incentivizing $\pi$ to achieve high proxy reward and stay close to the base policy, $\pi$ can achieve high true reward.
By scaling the RHS of (10) and removing terms that are constant with respect to $\pi$, we arrive at the following regularized policy optimization objective to avoid reward hacking:
@@
While Theorem 5.1 does not guarantee that the true reward can be increased by optimizing (5), it does at least allow us to provably avoid reward hacking. In Theorem A.3 in the appendix, we show that it is difficult to guarantee an improvement in true reward in general. However, if the lower bound on the RHS of (10) can be increased above zero—which can be tested empirically—then we know that the true reward has also increased. In our experiments in Section 6, we show that in many realistic environments it is possible to increase the true reward by optimizing (5).
Comparison to KL regularization in RLHF.  Currently, RLHF calls for the application of a KL penalty between the distributions of responses generated by the optimized policy and the SFT policy (Stiennon et al., 2020; Bai et al., 2022). In our setting, we can write this as
@@
Our regularized objective in (5) differs from the regularization applied in RLHF in two key ways: our approach uses occupancy measure (OM) divergence instead of action distribution (AD) divergence, and it employs $\chi^{2}$ divergence instead of KL divergence. In this section, we provide intuition for why choosing OM over AD divergence and $\chi^{2}$ over KL divergence are more effective for preventing reward hacking. Then, in Section 6, we empirically explore applying different types of regularization to prevent reward hacking in the five environments we study.
@@
Figure 4: Our theory also suggests reward hacking can more effectively be prevented by regularizing with $\chi^{2}$ divergence instead of KL divergence. This plot illustrates how $\chi^{2}$ regularization is more effective at preventing reward hacking in RLHF. Both divergences can be written as the expectation of a function $g\big(\log\big(\mu_{\pi}(s,a)/\mu_{{\pi_{\text{base}}}}(s,a)\big)\big)$ which increases the penalty on state-action pairs based on how far the log-ratio is from zero. The $g(\cdot)$ associated with KL divergence only increases slowly for large log-ratios, so policies trained with KL divergence may produce nonsensical text. In contrast, the $g(\cdot)$ for $\chi^{2}$ divergence increases exponentially, better constraining the LLM to produce text similar to the SFT policy.
Occupancy measure vs. action distribution regularization.  While AD regularization works well for RLHF, this may be because RLHF is essentially a contextual bandit problem, meaning that $\gamma=0$; in this case, OM and AD divergence are equivalent (see Appendix A.3). However, in other cases, AD divergence may not suffice to prevent reward hacking behavior. This is because, in longer-horizon environments, a small change in action distribution at a single state can lead to a much higher probability of reaching undesirable states. Figure 3 shows an example of this in the glucose monitoring environment: policies that are close in action distribution regularization produce vastly different patient glucose levels. Occupancy measure regularization avoids this issue by directly preventing the distribution of glucose levels from differing too much from the base policy. In Theorem A.5 in the appendix, we show that in general it is impossible to lower bound the improvement in true reward using almost any form of AD regularization, in contrast to our results on OM regularization.
$\chi^{2}$ vs. KL divergence.  Compared to KL divergence, $\chi^{2}$ divergence may be more effective for preventing reward hacking because it more strongly penalizes out-of-distribution state-action pairs. To illustrate this, we can write both divergences as expectations over functions of the log-ratio of the occupancy measures, which we denote $d(s,a)$:
@@
Table 1: We compare using various types of regularization to prevent reward hacking in the five environments form Figure 2. The median true reward and standard deviation across 5 random seeds is shown for the best regularization coefficient for each type of regularization. The bottom rows show results for the baselines: the base policy $\pi_{\text{base}}$}, a policy trained on the proxy reward without regularization (exhibiting reward hacking), and a policy trained on the true reward function (impossible in practice, but included as an upper bound on performance). We find that occupancy measure regularization consistently improves on action distribution regularization, and that $\chi^{2}$ divergence often outperforms KL divergence.
@@ 6 Experiments
We now show that our theoretical results—which suggest $\chi^{2}$ occupancy measure regularization can prevent reward hacking—translate to empirical success in realistic environments.
Occupancy measure regularization in practice.  Occupancy measure regularization is more difficult to implement in practice compared to action distribution regularization. While AD regularization can be added as a loss term to deep RL algorithms like proximal policy optimization (PPO) (Schulman et al., 2017), OM divergences cannot be calculated in closed form. Instead, we follow several previous works (e.g., Ho & Ermon 2016; Kang et al. 2018) and use a discriminator network to approximate OM divergences. Specifically, in Appendix B, we show the objective in (5) can be optimized via policy gradient with an adjusted reward function that depends on a discriminator $\hat{d}_{\phi}$:
@@
$\text{and}\qquad\phi$ $=\arg\min_{\phi^{\prime}}\;\mathbb{E}_{\mu_{\pi}}\Big[\,\log(1+e^{-\hat{d}_{\phi^{\prime}}(s,a)})\,\Big]+\,\mathbb{E}_{\mu_{\pi_{\text{base}}}}\Big[\,\log(1+e^{\hat{d}_{\phi^{\prime}}(s,a)})\,\Big].$ (8)
That is, optimizing a discriminator network $\hat{d}_{\phi}$ to minimize the given loss can be used to estimate and optimize $\chi^{2}$ OM divergence. We alternately train $\hat{d}_{\phi}$ via gradient descent on (8) and the policy $\pi$ via PPO based on the adjusted reward. We call this algorithm Occupancy-Regularized Policy Optimization (ORPO). See Appendix B for a full derivation of the approximations used in ORPO and Algorithm 1 for a formal description; we also describe how to regularize based on OM KL divergence.
@@
In Figure 5, we show the true reward achieved when training with each type of regularization across a range of $\lambda$ values. In addition to performing best with an optimal coefficient, $\chi^{2}$ and OM regularization seem to also perform well over a larger range of coefficients compared to AD regularization. In Appendix E, we present the full results of our experiments and ablations of ORPO.